### v1

# FOCUS: Fused Observation of Channels for Unveiling Spectra

Abstract
Hyperspectral imaging (HSI) captures hundreds of narrow, contiguous wavelength bands, making it a powerful tool in biology, agriculture, and environmental monitoring. However, interpreting Vision Transformers (ViTs) in this setting remains largely unexplored due to two key challenges: (1) existing saliency methods struggle to capture meaningful spectral cues, often collapsing attention onto the class token, and (2) full-spectrum ViTs are computationally prohibitive for interpretability, given the high-dimensional nature of HSI data. We present FOCUS, the first framework that enables reliable and efficient spatial–spectral interpretability for frozen ViTs. FOCUS introduces two core components: class-specific spectral prompts that guide attention toward semantically meaningful wavelength groups, and a learnable [SINK] token trained with an attraction loss to absorb noisy or redundant attention. Together, these designs make it possible to generate stable and interpretable 3D saliency maps and spectral importance curves in a single forward pass—without any gradient backpropagation or backbone modification. FOCUS improves band-level IoU by 15%, reduces attention collapse by over 40%, and produces saliency results that align closely with expert annotations. With less than 1% parameter overhead, our method makes high-resolution ViT interpretability practical for real-world hyperspectral applications—bridging a long-standing gap between black-box modeling and trustworthy HSI decision-making.
1 Introduction
[…]
Hyperspectral imaging (HSI) captures dense spectral signatures across hundreds of narrow, contiguous bands, enabling fine-grained analysis of physiological traits such as stress, disease, or pigment content. This spectral richness holds enormous promise for applications in biology, agriculture, and environmental monitoring. Recently, Vision Transformers (ViTs) have shown strong performance in hyperspectral recognition tasks, due to their ability to model long-range spatial–spectral dependencies.
However, such progress has not translated to interpretability. For real-world deployment, HSI models must answer not only “where” decisions are made (spatial saliency), but also “which wavelengths” are responsible—a need that current ViT interpretability tools fail to meet. Existing methods like Grad-CAM [32], LayerCAM [22], or Prompt-CAM [8] were designed for 3-channel RGB inputs, and cannot scale to the high-dimensional spectral input of HSI. They either ignore the spectral axis entirely or yield coarse spatial maps without wavelength reasoning—rendering them biologically uninformative. The core barrier is computational: unlike RGB, hyperspectral cubes often contain 200–300 channels, leading to input token sequences that are orders of magnitude larger. Since ViT attention scales quadratically with token count, spectral interpretability becomes intractable in both memory and runtime cost. In practice, this forces existing methods to avoid the full spectrum altogether—leaving a fundamental gap in saliency modeling. This is further exacerbated by the attention sink phenomenon [40], where attention collapses into dominant tokens such as [CLS]. In HSI, this issue worsens with spectral scale: the more spectral tokens, the more diluted and unstable the attention becomes. When ViTs fail to lock onto meaningful wavelength patterns, they default to class tokens, leading to spatial drift and spectral blindness (Fig. 1). For example, in plant disease detection, key reflectance signals often lie in narrow bands like the red-edge (700–750 nm). Existing methods fail to highlight such bands, producing diffuse or misleading saliency maps that misalign with expert knowledge and limit downstream utility.
We address these challenges by introducing FOCUS, the first framework to provide efficient and stable spatial–spectral interpretability for frozen ViTs on hyperspectral data. FOCUS introduces two lightweight but powerful innovations. First, we design class-specific spectral prompts that encode physiologically meaningful wavelength groups (e.g., VIS, red-edge, NIR), anchoring attention in the spectral domain. Second, we introduce a learnable [SINK] token trained with an attraction loss that explicitly absorbs redundant, low-relevance attention—transforming the attention sink from a failure mode into a controlled filtering mechanism.
With these components, FOCUS enables reliable interpretation of ViT-based HSI models with minimal overhead. It produces (i) a high-resolution 3D saliency cube across spatial and spectral dimensions, (ii) interpretable spectral attribution curves, and (iii) stable attention pathways without requiring test-time gradients or model modification. Extensive evaluation on two HSI benchmarks—including the new HyperLeaf2024 dataset—shows that FOCUS improves band-level IoU by 15%, reduces attention collapse by 40%, and generates domain-aligned explanations. These gains are achieved with under 1% additional parameters.
To summarize, our contributions are:
  • We are the first to identify and resolve a fundamental scalability barrier for ViT-based interpretability in hyperspectral imaging (HSI). Existing methods cannot scale to full-spectrum inputs due to the exponential growth in attention computation with increasing spectral channels. Our framework offers the first practical solution for interpretable ViTs under high-dimensional spectral conditions.
  • We introduce FOCUS, the first interpretability framework that jointly addresses spectral relevance and attention stability. By combining class-specific spectral prompts and a learnable [SINK] token with an attraction loss, FOCUS enables precise wavelength attribution while effectively filtering noisy or redundant attention—substantially mitigating the attention collapse phenomenon.
  • FOCUS is fully gradient-free, architecture-agnostic, and compatible with frozen ViTs. It generates high-resolution spatial–spectral explanations through a single forward pass, achieving state-of-the-art interpretability with less than 1% parameter overhead—making it practical for real-world deployment in biology and agriculture.
2 Related Work
2.1 Saliency Maps for Transformers
Initial CAM-based explanations emerged from gradient-centric methods, such as Grad-CAM [32], Eigen-CAM [27], Score-CAM [37], and XGrad-CAM [16]. As Transformers became prevalent, their noisy and layer-spanning gradient flows led to the development of attention-based alternatives, including AttnRollout [1], TokenGrad [46], and LRP-ViT [4], which propagate attention weights rather than backpropagate gradients. Smooth-CAM [28] further enhances robustness via integrated smoothing. More recent works [41, 42], such as Prompt-CAM [8], introduce learnable prompts to guide explanations, while DiffCAM [24] eliminates gradient reliance by comparing feature differences with reference distributions. Yet, despite their architectural innovations, all these methods operate solely in the spatial domain, lacking any mechanism to capture spectral relevance. This omission limits their applicability in domains like HSI, where spectral reasoning is foundational.
2.2 Hyperspectral Recognition and Explainability
Early approaches to HSI classification relied on 3D CNNs [25] or combinations of 2D spatial and 1D spectral convolutions [26], aiming to jointly encode local and spectral context. Recent methods leverage Transformer architectures—SST [7], HSViT [43], SpectralFormer [19], and HyperViT [31]—to capture long-range dependencies and achieve higher accuracy. However, these models typically expose only coarse spatial attention maps and offer limited interpretability in the spectral dimension. To address this, channel-level attribution methods such as SANet [13], SSA-CAM [33], and Hyperspectral Band [45] identify influential wavelengths but discard spatial granularity. Meanwhile, prompt-efficient tuning methods like HyperPrompt [18] and SPTrack-Prompt [17] improve transferability but lack explanatory depth. To date, no approach provides joint spatial–spectral saliency for frozen HSI ViTs—leaving a critical gap in understanding model decisions from both perspectives, largely due to the prohibitive computational cost of modeling dense spectral tokens in attention-based architectures.
2.3 Attention Sink in Vision Transformers
Vision Transformers (ViTs) frequently suffer from attention sink—a collapse of the attention distribution where a handful of tokens, most notably the class token, monopolise the receptive field, degrading robustness and rendering attribution maps unstable. Existing counter‑measures fall into two principal camps: Auxiliary‑token methods, such as SpecFormer, which flattens the attention spectrum by penalising large singular values during training [21]; Register‑ViT, which attaches register tokens that compete with the class token and absorb redundant activation [9]; and EDIT, which adds editing tokens encoding fine‑grained sub‑class structure that act as controllable sinks [15]; and Pruning and rescaling approaches, including TokenPruner, which progressively excises low‑saliency tokens at inference time [38, 39], Adaptive Token Sampler (ATS), which learns an importance mask that discards uninformative heads on‑the‑fly [14], and LayerScale, which mitigates early‑layer domination by rescaling residual branches with learnable coefficients [35]. Although effective, these techniques typically require full retraining and do not provide a quantitative gauge of collapse mitigation. By appending a single [SINK] token and a lightweight attraction loss that siphons non‑discriminative attention into an explicit, interpretable reservoir, our method serves as a drop‑in module for frozen backbones and, for the first time, enables direct measurement of collapse mitigation through two metrics: Sink‑Rate—the proportion of attention mass absorbed by [SINK]—and Sink‑Consistency—the inter‑layer stability of saliency maps.
3 Method
3.1 Overview
Given a hyperspectral image cube $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$ with $C$ spectral bands and spatial resolution $H\times W$, our goal is twofold: (i) accurately predict a target outcome, such as plant genotype, disease status, or physiological traits, and (ii) generate an interpretable joint spatial–spectral saliency map, clearly indicating both where and which wavelengths influence model decisions.
To achieve these objectives, we extend a frozen Vision Transformer (ViT) [12] backbone by introducing three novel and intuitive modules:
  1. 1.
    Class-specific spectral prompts, explicitly guiding wavelength-aware attention (Section 3.2);
  2. 2.
    A learnable [SINK] token with an associated attraction loss to convert attention collapse [40] into a controllable noise-filtering mechanism (Section 3.3);
  3. 3.
    An attention aggregation routine, forming a unified three-dimensional saliency cube from prompt-to-patch attention (Section 3.4).
Only the spectral prompts, [SINK] token, and a lightweight spectral adapter are trainable, collectively adding fewer than 1% parameters to the ViT. Importantly, inference is fully gradient-free, as illustrated in Figure 2.
Figure 2: Overview of FOCUS. A hyperspectral image is embedded and processed by a frozen ViT. FOCUS injects class-specific spectral prompts, each corresponding to meaningful spectral groups (e.g., VIS, red-edge, NIR), and introduces a dedicated [SINK] token to absorb redundant attention. Saliency is computed via prompt-to-patch attention, yielding a 3D attribution cube $\mathbf{T}\in\mathbb{R}^{H\times W\times C}$ for spatial–spectral interpretation.
3.2 Class-Specific Spectral Prompts
We partition the $C$ spectral bands into $G$ non-overlapping groups $\{\mathcal{B}_{g}\}_{g=1}^{G}$, guided by domain knowledge [20], such as visible (VIS: 400–700 nm), red-edge (700–750 nm), NIR (750–1100 nm), and SWIR (1100 nm). Let $C_{g}=|\mathcal{B}_{g}|$ denote the number of bands in group $g$.
For each class $c\in\{1,\dots,K\}$ and spectral group $g$, we introduce a learnable spectral prompt token $\mathbf{p}_{c,g}\in\mathbb{R}^{d}$. The input sequence to the first ViT layer is:
$\mathbf{Z}^{(0)}=\left[\underbrace{\mathbf{s}}_{\text{\scriptsize[SINK]}},\;\underbrace{\mathbf{p}_{c,1},\dots,\mathbf{p}_{c,G}}_{\text{spectral prompts}},\;\underbrace{\mathbf{x}_{1},\dots,\mathbf{x}_{N}}_{\text{patch tokens}}\right],$ (1)
where $N=H\cdot W$ is the number of spatial patches, and each spectral band is embedded using a depth-wise adapter $\boldsymbol{\Phi}\in\mathbb{R}^{C\times d}$ before patchification.
Each spectral prompt token $\mathbf{p}_{c,g}$ serves as a “spectral anchor,” exclusively interacting with patches from its assigned spectral group $\mathcal{B}_{g}$. This design ensures sharper attention and enhanced spectral interpretability compared to generic prompts [8, 18].
3.3 Sink-Aware Attention Routing
At transformer layer $\ell$ and head $h$, attention is computed via standard scaled dot-product:
$\mathbf{A}^{\ell,h}=\operatorname{softmax}\left(\frac{\mathbf{Q}^{\ell,h}\mathbf{K}^{\ell,h\top}}{\sqrt{d}}\right),$ (2)
where $\mathbf{Q},\mathbf{K}\in\mathbb{R}^{(1+G+N)\times d}$, and the [SINK] token index is set as $k_{\mathrm{sink}}=1$.
We identify a subset of low-informativeness heads $\mathcal{H}_{\text{aux}}$ [34, 10] to serve as noise absorbers. We introduce a lightweight attraction loss encouraging these heads to route maximal attention to the [SINK] token:
$\mathcal{L}_{\text{sink}}=-\lambda\sum_{\ell}\frac{1}{|\mathcal{H}_{\text{aux}}|}\sum_{h\in\mathcal{H}_{\text{aux}}}\operatorname{mean}_{i}\left[\mathbf{A}^{\ell,h}_{i,k_{\mathrm{sink}}}\right],$ (3)
where $\lambda$ is a small scalar (e.g., $10^{-3}$). This strategy mitigates attention collapse, enabling discriminative heads to better capture spatial–spectral signals.
3.4 3-D Prompt-CAM Generation
After the final transformer layer $L$, we aggregate prompt-to-patch attention from discriminative heads =$\mathcal{H}_{\star}=\{1,\dots,H\}\setminus\mathcal{H}_{\text{aux}}$ into a spatial–spectral saliency cube:
$T_{x,y,\lambda}=\frac{1}{|\mathcal{H}_{\star}|}\sum_{h\in\mathcal{H}_{\star}}\mathbf{A}^{L,h}_{q=\mathbf{p}_{c,g(\lambda)},\,k=\mathbf{x}_{x,y,\lambda}},$ (4)
where $g(\lambda)$ denotes the spectral group index of band $\lambda$. The resulting cube $\mathbf{T}\in\mathbb{R}^{H\times W\times C}$ simultaneously encodes spatial and spectral importance.
To facilitate intuitive interpretation, we define a spatial heatmap $M_{x,y}$ and a spectral importance curve $B_{\lambda}$ as:
$M_{x,y}=\sum_{\lambda}T_{x,y,\lambda},\quad B_{\lambda}=\max_{x,y}T_{x,y,\lambda}.$ (5)
Both visualizations are obtained efficiently in a single forward pass, without gradients.
3.5 Training Objective and Complexity
Our complete training objective combines task-specific loss $\mathcal{L}_{\text{task}}$ (cross-entropy or MSE) and the attention attraction loss $\mathcal{L}_{\text{sink}}$:
$\mathcal{L}=\mathcal{L}_{\text{task}}+\mathcal{L}_{\text{sink}}.$ (6)
During training, the ViT backbone remains frozen, while only (i) the spectral prompts ($K\times G\times d$), (ii) the [SINK] token ($d$), and (iii) the depth-wise spectral adapter ($C\times d$) are optimized. Collectively, these parameters add fewer than 1% overhead relative to the original ViT-B/16 model (0.8M vs. 86M parameters), introducing only a marginal computational cost (approximately 2.1% additional FLOPs).
Our method requires no gradient computation during inference. A single forward pass simultaneously provides accurate predictions, a comprehensive 3D saliency cube, a clear spatial heatmap, and a precise spectral relevance curve, directly supporting downstream analysis and biological interpretation.
4 Experiments
4.1 Experimental Setup
Datasets
We evaluate FOCUS on two representative hyperspectral leaf datasets to comprehensively evaluate its performance in both classification and regression settings. The first, HyperLeaf 2024 [23], comprises 2,410 VNIR–SWIR hyperspectral image cubes, each with a spatial resolution of $256{\times}256$. Each sample is annotated with both genotype labels (4 classes) and continuous physiological indicators such as chlorophyll ab and water content, enabling the simultaneous study of discrete classification and continuous trait regression. The second dataset, Tomato BLS [47], contains 3,120 hyperspectral images of tomato leaves, labeled as either healthy or infected with bacterial leaf spot (BLS). This binary disease classification task is particularly suitable for evaluating the precision of saliency methods in localizing stress-affected regions and detecting spectrally discriminative cues under real-world noise and biological variation.
Evaluation Metrics
We adopt a suite of established metrics that collectively capture classification accuracy, spectral attribution fidelity, and spatial alignment with expert annotations. Classification Accuracy (Acc) measures standard top-1 prediction correctness. Band-level IoU (BIO@5), introduced by SANet [13], quantifies spectral interpretability by computing Intersection-over-Union over the top-5 most influential wavelengths. Spatial IoU, following the DiffCAM protocol [24], evaluates overlap between predicted saliency regions and ground-truth lesion masks. Lastly, AUPRC (Area Under Precision–Recall Curve) is used to evaluate fine-grained spatial localization quality, especially in class-imbalanced disease identification scenarios. This evaluation protocol aligns with prevailing practices in the hyperspectral interpretability literature [13, 24, 8]. All results are averaged over three random dataset splits to ensure robustness.
Baselines
We compare FOCUS against a comprehensive suite of 18 state-of-the-art explainability methods spanning four major categories. Classical CAM-based techniques include Grad-CAM [32], Grad-CAM++ [3], Score-CAM [37], Eigen-CAM [27], XGrad-CAM [16], Layer-CAM [22], and Ablation-CAM [11], all of which rely on gradient-based activation visualization. Transformer-specific interpretability methods include AttnRollout [1], TokenGrad [46], LRP-ViT [4], Prompt-CAM [8], and DiffCAM [24], which exploit attention flow and token saliency for ViT explanations. Prototype- and data-driven models such as ProtoPNet [6], ProtoPFormer [44], INTR [29], and RISE [30] provide example-based or perturbation-based interpretability. Finally, we also evaluate two hyperspectral-specific approaches—SANet [13] and SSA-CAM [33]—that directly model spectral importance but typically lack spatial resolution. For fair comparison, all baselines are re-implemented using a common frozen ViT-B/16 (DINOv2) backbone and applied to hyperspectral cubes preprocessed under the same pipeline. Evaluation protocols follow consistent saliency metrics across spatial, spectral, and joint dimensions, ensuring a level playing field.
Implementation Details
We freeze the pretrained ViT-B/16 backbone (DINOv2) during all experiments, training only the proposed spectral prompts, the [SINK] token, and a lightweight $1\times 1$ spectral adapter. We employ the AdamW optimizer with parameters ($\beta_{1}=0.9$, $\beta_{2}=0.95$), learning rate $5\times 10^{-4}$, and weight decay $10^{-2}$, training for a total of 100 epochs. The number of spectral prompt groups ($G$) is set to 10 based on ablation results, and the sink attraction loss weight ($\lambda$) is set to $10^{-3}$.
Figure 3: Band-wise interpretability across spectral groups. Compared to baseline methods (Grad-CAM and Prompt-CAM), FOCUS produces significantly clearer and biologically grounded saliency patterns, especially in the red-edge and shortwave infrared (SWIR) regions (columns 4–5). The red-edge band (700–750 nm) is known to reflect changes in chlorophyll concentration and leaf pigment activity, critical for identifying early-stage stress and disease. The SWIR band (around 2,100 nm) is highly sensitive to water content and cellular structure, offering key cues for detecting drought or tissue damage. Our method highlights these regions with greater clarity and spatial precision, aligning closely with expert physiological annotations and enhancing model transparency in plant health assessment.
4.2 Main Results
Table 1 presents the main quantitative results of FOCUS and competitive baselines across two challenging hyperspectral benchmarks: HyperLeaf 2024 and Tomato BLS. For each dataset, we report the most informative and widely-used metrics: classification accuracy (Acc), band-level interpretability (BIO@5), spatial localization quality (IoU), and overall explanation precision (AUPRC). All methods share an identical ViT-B/16 (DINOv2) backbone and consistent training/evaluation protocols. On both datasets, FOCUS achieves the best performance in all reported metrics, demonstrating not only strong predictive ability, but also superior spatial and spectral interpretability. Notably, on HyperLeaf 2024, FOCUS improves BIO@5 and IoU by over 15% compared to the best transformer and HSI-specific baselines. On Tomato BLS, our approach yields sharper lesion localization and more accurate band attribution, further highlighting its generalizability and robustness.
Table 1: Main results across multiple hyperspectral benchmarks. For each dataset, we report classification accuracy (Acc), band-level interpretability (BIO@5), spatial localization (IoU), and explanation precision (AUPRC). Best results are in bold. FOCUS consistently leads on both predictive and interpretability metrics.
Method HyperLeaf 2024 Tomato BLS
Acc $\uparrow$ BIO@5 $\uparrow$ IoU $\uparrow$ AUPRC $\uparrow$ Acc $\uparrow$ BIO@5 $\uparrow$ IoU $\uparrow$ AUPRC $\uparrow$
Grad-CAM [32][ICCV’17] 82.1 0.33 0.41 0.53 79.3 0.28 0.39 0.51
SANet [13][TIP’20] 83.0 0.36 0.43 0.56 80.5 0.29 0.41 0.52
Prompt-CAM [8][CVPR’25] 84.5 0.38 0.46 0.59 81.2 0.34 0.43 0.58
DiffCAM [24][CVPR’25] 85.3 0.41 0.49 0.62 82.4 0.36 0.47 0.61
FOCUS (Ours) 87.0 0.56 0.58 0.72 84.6 0.47 0.54 0.70
Table 2: Efficiency and inference overhead. FOCUS offers minimal overhead and needs no gradients.
Method Params FLOPs Grad Latency
(%) (%) (ms)
Grad-CAM [32][ICCV’17] 0.0 12.5 11.2
SANet [13][TIP’20] 2.1 8.2 10.5
Prompt-CAM [8][CVPR’25] 1.3 5.6 8.6
DiffCAM [24][CVPR’25] 0.5 4.8 7.9
FOCUS (Ours) 0.9 2.1 6.8
Efficiency and Inference Overhead.
As shown in Table 2, FOCUS incurs less than 1% parameter overhead and achieves the lowest FLOPs and runtime latency among all baselines. Importantly, it operates in a fully gradient-free manner at inference, which sharply reduces deployment cost compared to gradient-based methods (e.g., Grad-CAM [32], SANet [13]). These findings are consistent with recent work [8, 18], highlighting that prompt-based and gradient-free explainers enable practical, real-time visual interpretation for large transformer models.
4.3 Ablation Study
Table 3: Spectral prompt granularity. Moderate group count ($G=10$) achieves best interpretability without excess overhead.
#Groups $G$ Acc $\uparrow$ BIO@5 $\uparrow$ $\rho$uparrow Params (%)
2 91.1 0.48 0.63 0.4
5 92.6 0.52 0.68 0.6
10 93.7 0.56 0.72 0.9
20 92.9 0.54 0.70 1.6
Table 4: Sink token and loss. Both accuracy and explanation stability improve with explicit sink routing.
Config Acc $\uparrow$ BIO@5 $\uparrow$ Sink Rate $\downarrow$ Sink Consist. $\uparrow$
w/o Sink 93.5 0.51 0.38 0.64
+ Sink, no loss 93.4 0.53 0.31 0.69
+ Sink & Loss 93.7 0.56 0.22 0.75
Spectral Prompt Granularity.
Table 3 shows that increasing the number of spectral groups ($G$) improves both accuracy and interpretability, but only up to a moderate value ($G=10$), beyond which gains saturate or slightly decline. This echoes prior findings [18, 13] that over-partitioning can dilute spectral focus and increase model overhead.
Sink Control Mechanism.
Table 4 demonstrates that introducing a dedicated [SINK] token and the associated attraction loss yields clear improvements in both interpretability and explanation stability, reducing spurious attention accumulation. This is consistent with recent work on attention regularization and prompt-based filtering in transformers [40, 10, 18].
4.4 Qualitative Analysis
We present visual comparisons to qualitatively evaluate the interpretability and spectral coherence of FOCUS. FOCUS produces cleaner and more localized saliency patterns than existing methods. Figure 3 illustrates that Grad-CAM tends to generate weak, spatially diffuse saliency, offering limited interpretive value under challenging spectral conditions. Prompt-CAM shows improved contrast but still lacks fine-grained localization in biologically informative regions.
In contrast, FOCUS produces sharp, consistent attention maps that tightly align with known physiological markers—specifically, stress-induced lesions in the spatial domain and well-defined spectral responses at the red-edge (750 nm) and SWIR ( 2,100 nm) bands, which are strongly linked to chlorophyll degradation and protein absorption in plant tissues. This highlights the biological validity and interpretability of our approach. In Figure 4, we observe that FOCUS is robust across replicates: attention consistently focuses on biochemically informative areas, whereas Prompt-CAM shows unstable activation and Grad-CAM fails to distinguish background noise. This suggests that our spectral prompts and sink routing modules provide more stable and reproducible attention allocation. Finally, Figure 6 shows band-wise heatmaps under different spectral groupings. FOCUS not only identifies where important features lie spatially, but also which specific bands matter—e.g., sharp peaks in the red-edge and protein-sensitive SWIR bands—demonstrating fine-grained spatial–spectral attribution. By contrast, baseline methods either collapse saliency across bands or highlight spectrally irrelevant regions, offering little insight to domain experts. Together, these qualitative findings reinforce the quantitative results and validate FOCUS as a reliable tool for interpretable hyperspectral vision tasks.
Figure 4: Saliency maps across spatial and spectral dimensions. Compared to Grad-CAM and Prompt-CAM, FOCUS exhibits both sharper spatial saliency and cleaner spectral separation, especially at red-edge (750 nm) and SWIR (2,100 nm) bands.
4.5 Analysis of Attention Sink Behavior
To elucidate how attention collapse manifests in ViT models on hyperspectral data, we analyze the distribution of attention mass routed to the [CLS] and [SINK] tokens across layers and heads. Figure 7 visualizes the per-head attention entropy and sink absorption rate under three configurations: (i) vanilla ViT, (ii) ViT with [SINK], and (iii) ViT with both [SINK] and the proposed attraction loss. Without explicit sink routing, over 30% of attention mass consistently accumulates on the first token, echoing the “attention-sink” phenomenon observed in recent ViT literature [40, 48]. By introducing our structured sink mechanism, this rate drops sharply, while predictive performance remains unchanged (see Table 4). This demonstrates that FOCUS does not merely regularize attention, but actively enables stable, trustworthy, and interpretable attention routing for long-sequence ViTs. Similar findings have been reported in recent XAI studies that combine architectural innovations with explicit routing or gating for robust interpretability [8, 18].
4.6 Robustness Under Spectral Noise
Hyperspectral imaging is inherently susceptible to noise at high wavelengths (especially SWIR $>$2200 nm) due to sensor artifacts, water absorption, or atmospheric effects. This not only reduces model reliability, but can also destabilize saliency explanations as transformers overfit to spurious spectral patterns. Empirically, we observe that vanilla ViT-based saliency methods (e.g., Grad-CAM [32], Prompt-CAM [8]) often produce volatile or physiologically irrelevant attributions in these noisy bands—a phenomenon also documented in [48, 5].
In contrast, our sink mechanism acts as a noise valve: unstable heads route high-entropy responses to [SINK], selectively suppressing noise-induced activations. This yields more robust and biologically meaningful explanations, without additional model tuning. Our results resonate with the “attention debiasing” literature, which finds that targeted attention steering—whether by loss design [5] or prompt gating [2]—improves model resilience to common corruptions. Uniquely, FOCUS achieves these gains with less than 1% additional parameters and without retraining the ViT backbone.
4.7 Computational Efficiency Comparison
Interpretability in hyperspectral vision must be efficient to support deployment in real-world agricultural and biological scenarios with limited resources. However, standard ViTs scale poorly for hyperspectral cubes: a single image may contain 200+ spectral bands, leading to quadratic growth in memory and attention cost. Direct inference with a vanilla ViT requires over 14,GB of GPU memory. Recent approaches like D-CHAG [36] use spectral aggregation to reduce computation, yet still consume 8.4,GB and require model modifications and retraining. In contrast, FOCUS introduces no backbone changes or gradients. Through lightweight spectral prompts and a learnable [SINK] token, it requires only 1.9,GB of memory and operates in a single forward pass. As shown in Figure 5, FOCUS reduces memory cost by over 85% while delivering fine-grained spatial–spectral explanations. This efficiency makes our method the first practically deployable interpretability solution for ViTs in high-dimensional hyperspectral vision.
Figure 5: Inference efficiency comparison. FOCUS achieves over 95% runtime and 85% memory reduction compared to vanilla ViT and D-CHAG [36], while providing complete spatial–spectral interpretability.
4.8 Limitations
Despite its strong interpretability and low overhead, FOCUS has two current limitations. First, the number of spectral prompts grows linearly with the number of target classes ($K$), which may impact scalability in large-scale recognition tasks. Future work may explore shared or hierarchical prompting strategies to reduce prompt complexity. Second, under severe spectral noise (e.g., beyond 2200 nm in SWIR), the [SINK] token may over-filter faint yet informative signals. Integrating adaptive routing or uncertainty-aware mechanisms could offer better trade-offs between suppression and retention.
Figure 6: Band-wise saliency comparison at key wavelengths. Visualization of saliency maps at 750 nm (red-edge), 1100 nm (NIR), and 2100 nm (SWIR). Compared to Grad-CAM and Prompt-CAM, our method produces clearer and more concentrated saliency, particularly at red-edge and SWIR bands, which are known to be physiologically relevant for leaf health. The colorbar reflects normalized attention intensities.
Figure 7: Layer-wise sink attention mass. Comparison of attention mass routed to the [SINK] token across transformer layers. Our full method (+Sink+$\mathcal{L}_{\text{sink}}$) most effectively drains non-discriminative attention, reducing accumulation over depth.
5 Conclusion
We introduced FOCUS, a lightweight and gradient-free interpretability framework tailored for hyperspectral vision transformers. By combining class-specific spectral prompts with a learnable [SINK] token and a simple attraction loss, FOCUS produces stable and fine-grained spatial–spectral saliency maps. Despite adding less than 1% additional parameters, it achieves state-of-the-art interpretability performance—improving spectral IoU by 15% and enhancing attribution consistency across layers. Our work addresses a fundamental scalability barrier in hyperspectral interpretability, enabling trustworthy and deployable explanations for high-dimensional spectral data. We believe FOCUS lays a practical foundation for interpretable AI in spectral-intensive domains. In future work, we aim to explore adaptive prompt sharing and hierarchical strategies to further scale FOCUS to long-sequence spectral tasks.
Acknowledgments
This manuscript was authored by Oak Ridge National Laboratory (ORNL), operated by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S. Department of Energy. Any subjective views or opinions expressed in this paper do not necessarily represent those of the U.S. Department of Energy or the United States Government.